OpenStudy (anonymous):

If tan(x+50)=cotx a value of x is... Explain please

OpenStudy (usukidoll):

isn't there an identity formula for tangent and cotangent you can use?!

OpenStudy (anonymous):

Convert into sin and cos first.

OpenStudy (anonymous):

Then you will end up with an angle sum formula.

OpenStudy (anonymous):

I think you'll need to use \[ \sin(a+b) = \sin(a)\cos(b) - \sin(b)\cos(a) \]

OpenStudy (***[isuru]***):

wio... wouldnt it be much easy if u write cot x = tan ( pi/2 -x) then u have the expression tan (x + 50) = tan ( pi/2 - x) so u can solve it directly.... or u can use the formula and find whole set of answers ? :)